Raisins cost $2 per pound and nuts costs $5 per pound. How many pounds of each should you use to make a 30-lb mixture that costs $4 per pound?
Answer by lwsshak3(11628)

Raisins cost $2 per pound and nuts costs $5 per pound. How many pounds of each should you use to make a 30-lb mixture that costs $4 per pound?
**
let x=amt of raisins to mix
30-x=amt of nuts to mix
..
2x+5(30-x)=4*30
2x+150-5x=120
3x=30
x=10
30-x=20
amt of raisins to mix=10 lbs
amt of nuts to mix=20 lbs